Delph Company uses a job-order costing system and has two manufacturing departments—Molding and Fabrication. The company provided the following estimates at the beginning of the year:

  

Molding Fabrication Total
Machine-hours 31,000 41,000 72,000
Fixed manufacturing overhead costs $ 790,000 $ 290,000 $ 1,080,000
Variable manufacturing overhead cost per machine-hour $ 5.70 $ 5.70

  

During the year, the company had no beginning or ending inventories and it started, completed, and sold only two jobs—Job D-70 and Job C-200. It provided the following information related to those two jobs:

  

Job D-70: Molding Fabrication Total
Direct materials cost $ 375,000 $ 325,000 $ 700,000
Direct labor cost $ 250,000 $ 160,000 $ 410,000
Machine-hours 24,000 7,000 31,000

  

Job C-200: Molding Fabrication Total
Direct materials cost $ 260,000 $ 230,000 $ 490,000
Direct labor cost $ 110,000 $ 250,000 $ 360,000
Machine-hours 7,000 34,000 41,000

Delph had no underapplied or overapplied manufacturing overhead during the year.

2. Assume Delph uses departmental predetermined overhead rates based on machine-hours.

a. Compute the departmental predetermined overhead rates.

b. Compute the total manufacturing cost assigned to Job D-70 and Job C-200.

c. If Delph establishes bid prices that are 150% of total manufacturing costs, what bid prices would it have established for Job D-70 and Job C-200?

d. What is Delph’s cost of goods sold for the year?
Answer #1

2a) Departmental overhead rate :

Molding department = (790000/31000)+5.7 = 31.18 per hour

Fabrication department = (290000/41000)+5.7 = 12.77 per hour

2b) Total manufacturing cost

Job D-70 Job C-200
Direct material 700000 490000
Direct labor 410000 360000
Molding department 24000*31.18 = 748320 218260
Fabrication department 89390 434180
Total manufacturing cost 1947710 1502440

2c) Bid price

Job D-70 = 1947710*1.5 = 2921565

Job C-200 = 1502440*1.5 = 2253660

2d) Cost of goods sold = 1947710+1502440 = 3450150